# The Living Forecast: Evolving Day-Ahead Predictions into Intraday Reality

### v1

@@ Abstract
Accurate intraday forecasts are essential for power system operations, complementing day-ahead forecasts that gradually lose relevance as new information becomes available. This paper introduces a Bayesian updating mechanism that converts fully probabilistic day-ahead forecasts into intraday forecasts without retraining or re-inference. The approach conditions the Gaussian mixture output of a conditional variational autoencoder-based forecaster on observed measurements, yielding an updated distribution for the remaining horizon that preserves its probabilistic structure. This enables consistent point, quantile, and ensemble forecasts while remaining computationally efficient and suitable for real-time applications. Experiments on household electricity consumption and photovoltaic generation datasets demonstrate that the proposed method improves forecast accuracy up to 25% across likelihood-, sample-, quantile-, and point-based metrics. The largest gains occur in time steps with strong temporal correlation to observed data, and the use of pattern dictionary–based covariance structures further enhances performance. The results highlight a theoretically grounded framework for intraday forecasting in modern power systems.
@@ I Introduction
Figure 1: Overview of the proposed Bayesian intraday update mechanism. A fully probabilistic day-ahead forecast, represented as a Gaussian mixture model, is obtained from a conditional variational autoencoder (top). Given partial intraday observations, the model applies a closed-form Bayesian update to adjust the remaining-horizon distribution (bottom), enabling consistent quantile, ensemble, and point forecasts without re-training or re-inference.
@@
A variety of methods have been explored for this purpose, but each comes with important limitations. One option is to train separate models for intraday forecasting, conditioned on partial observations [3], but this increases training and maintenance effort and may be inconsistent with the original day-ahead forecast. Sliding-window-based autoregressive (re-)training adapts to recent conditions, yet the required amount of data gets larger with the time resolution and it may underperform for longer horizons. Heuristic updates, such as bias correction or smoothing [1], are simple and fast but limited to marginal adjustments without a coherent probabilistic structure. Copula-based updates adjust temporal dependencies while preserving marginals [9]; however, separating marginal and dependency modelling can introduce inconsistencies. Overall, these methods can improve short-term accuracy but often demand significant resources or sacrifice probabilistic coherence.
In this work, we propose a simple and theoretically grounded approach for intraday forecast updating that directly builds on the existing day-ahead probabilistic forecast. Rather than retraining models, modifying architectures, or generating new scenarios from scratch, our method applies a Bayesian update to the original forecast distribution, conditioned on newly available intraday observations. The proposed update process is illustrated in Fig. 1: starting from the fully probabilistic day-ahead forecast distribution, the method conditions on the observed portion of the day to produce an updated distribution for the remaining hours. The underlying day-ahead forecast is modelled using a conditional variational autoencoder with a Gaussian mixture output, enabling rich and flexible representation of uncertainty across time. This structure enables the update to be performed directly in the distribution space, without requiring retraining or re-running the neural network. As a result, the updated forecast remains fully probabilistic, supports sampling and quantile extraction, and preserves consistency with the original model. The approach is computationally efficient and designed with practical applicability in mind, making it a promising candidate for integration into real-time forecasting pipelines.
@@ III-A Day-ahead forecasting model
$p_{\psi}(\mathbf{x}|\mathbf{z},\mathbf{c})=\mathcal{N}(\mathbf{x};\mu=f_{\psi}^{\mu}(\mathbf{z,c}),\Sigma=f_{\psi}^{\Sigma}(\mathbf{z,c}))$ (3)
@@ III-B Intraday updating mechanism
Once the day-ahead forecast $\tilde{p}_{\psi}(\mathbf{x}|\mathbf{c})$ is obtained, we aim to reuse it for intraday forecasting without any re-inference [^1] We refer to re-inference as running another forward pass on the neural network. or re-training. That is, we adapt the GMM output of the CVAE directly, as illustrated in Fig. 1.
@@ IV-B3 Quantile-based
We use the same ensembles to form quantile forecasts $y_{t,n,i}^{T^{\prime}:}$ representing the $q_{i}$-th quantile of the ensemble set $\{x_{t,n,s}^{T^{\prime}:}\}_{s=1}^{S}$. Here, $\{q_{i}\}_{i=1}^{Q}\subset[0,1]^{Q}$ is the collection of $Q$ quantile levels with $q_{i}<q_{i+1},~\forall i$. After obtaining the quantile forecasts for each test day, the pinball loss function can be used as the approximation of the well-known continuous ranked probability score (CRPS): [15]
@@ IV-C Baselines
The performance of the proposed intraday forecaster is heavily tied to the modelling accuracy of the day-ahead forecasting distribution. To isolate the effects of the updating mechanism on the day-ahead forecaster, we define a baseline model and a synthetic test set.
@@ V Results and Discussion
Finally, we compare performance over real and synthetic test sets in Fig. 6 and Fig. 7. For the Gipuzkoa SM dataset, performance traces are largely consistent with each other, except for the NLL metric, which suggests that the real test set contains more typical, and therefore easier-to-predict, samples. For the Utrecht PV dataset, on the other hand, the model performs better on the synthetic test set, as expected. Despite this discrepancy, we observe that the update mechanism still provides a performance gain on the real test set in all metrics.
@@ VI Conclusion
The experimental results on both electricity consumption and PV generation datasets confirm that the updating mechanism consistently improves forecast accuracy across likelihood-, sample-, quantile-, and point-based metrics. The improvements are most noticeable for time steps with strong temporal correlation to the observed data, including correlations that extend over longer ranges. This highlights the method’s ability to effectively incorporate real-time information. Moreover, the use of pattern dictionary–based covariance structures enhances both the baseline modelling capability and the quality of intraday updates, underlining the importance of explicit correlation modelling.

### v2

@@ Abstract
Accurate intraday forecasts are essential for power system operations, complementing day-ahead forecasts that gradually lose relevance as new information becomes available. This paper introduces a Bayesian updating mechanism that converts fully probabilistic day-ahead forecasts into intraday forecasts without retraining or re-inference. The approach conditions the Gaussian mixture output of a conditional variational autoencoder-based forecaster on observed measurements, yielding an updated distribution for the remaining horizon that preserves its probabilistic structure. This enables consistent point, quantile, and ensemble forecasts while remaining computationally efficient and suitable for real-time applications. Experiments on household electricity consumption and photovoltaic generation datasets demonstrate that the proposed method improves forecast accuracy up to 25% across likelihood-, sample-, quantile-, and point-based metrics. The largest gains occur in time steps with strong temporal correlation to observed data, and the use of pattern dictionary-based covariance structures further enhances performance. The results highlight a theoretically grounded framework for intraday forecasting in modern power systems.
@@ I Introduction
Figure 1: Overview of the proposed Bayesian intraday update mechanism. A fully probabilistic day-ahead forecast, represented as a Gaussian mixture model, is obtained from a conditional variational autoencoder (top). Given partial intraday observations, the model applies a closed-form Bayesian update to adjust the remaining-horizon distribution (bottom), enabling consistent quantile, ensemble, and point forecasts without retraining or re-inference.
@@
A variety of methods have been explored for intraday forecast refinement, yet each presents important limitations. One common approach is to train separate models for intraday forecasting conditioned on partial observations [14]. While effective, this strategy increases training and maintenance effort and may lead to inconsistencies with the original day-ahead forecast. Adaptive schemes based on sliding-window autoregressive training update predictions using recent data, but their data requirements grow with temporal resolution, and they may struggle to preserve performance over longer horizons. Heuristic corrections, such as bias adjustment or smoothing [7], are computationally efficient but operate primarily at the marginal level, lacking a coherent probabilistic structure. Copula-based approaches explicitly model temporal dependencies while preserving marginal distributions [1]. However, separating marginal and dependency modelling can introduce structural inconsistencies.
Recent probabilistic forecasting studies further emphasise the importance of joint uncertainty modelling in intraday power system applications. Ensemble-based mixture density networks have been applied to intraday PV forecasting to provide flexible distributional outputs [5], and Bayesian hierarchical frameworks have been proposed for continuous intraday electricity markets [8]. Multivariate update models for renewable infeed explicitly account for temporal correlations [7], while hybrid approaches combining Kalman post-processing with machine learning aim to refine intraday uncertainty estimates sequentially [12]. Advanced deep learning architectures integrating transformer models with Gaussian process approximations further demonstrate the potential of expressive probabilistic modelling [16].
Despite these advances, most existing approaches either require retraining, repeated neural network inference, or architectural modifications to incorporate new observations during the day. As a result, maintaining structural consistency with the original day-ahead joint distribution while ensuring computational scalability remains challenging.
In this work, we propose a simple and theoretically grounded approach for intraday forecast updating that directly builds on the existing day-ahead probabilistic forecast. Rather than retraining models, modifying architectures, or generating new scenarios from scratch, our method applies a Bayesian update to the original forecast distribution, conditioned on newly available intraday observations. The proposed update process is illustrated in Fig. 1: starting from the fully probabilistic day-ahead forecast distribution, the method conditions on the observed portion of the day to produce an updated distribution for the remaining hours. The underlying day-ahead forecast is modelled using a conditional variational autoencoder with a Gaussian mixture output, enabling rich and flexible representation of uncertainty across time. This structure enables the update to be performed directly in the distribution space, without requiring retraining or re-running the neural network. As a result, the updated forecast remains fully probabilistic, supports sampling and quantile extraction, and preserves consistency with the original model. The approach avoids retraining and additional neural network inference during intraday operation, relying instead on closed-form linear algebraic updates whose complexity scales with the number of observed time steps. This makes it suitable for integration into real-time forecasting pipelines in practical power system applications.
@@ III-A Day-ahead forecasting model
$p_{\psi}(\mathbf{x}|\mathbf{z}_{k},\mathbf{c})=\mathcal{N}(\mathbf{x};\mu=f_{\psi}^{\mu}(\mathbf{z}_{k},\mathbf{c}),\Sigma=f_{\psi}^{\Sigma}(\mathbf{z}_{k},\mathbf{c}))$ (3)
@@ III-B Intraday updating mechanism
Once the day-ahead forecast $\tilde{p}_{\psi}(\mathbf{x}|\mathbf{c})$ is obtained, we aim to reuse it for intraday forecasting without any re-inference [^1] We refer to re-inference as running another forward pass on the neural network. or retraining. That is, we adapt the GMM output of the CVAE directly, as illustrated in Fig. 1.
@@
Computational considerations
The intraday update requires only closed-form linear-algebraic operations and no additional neural network inference. For each mixture component, the dominant cost stems from inverting $\Sigma_{k}^{:T^{\prime}}\in\mathbb{R}^{T^{\prime}\times T^{\prime}}$, resulting in a per-component complexity of $\mathcal{O}(T^{\prime 3})$, where $T^{\prime}$ denotes the number of observed time steps. The remaining matrix multiplications scale as $\mathcal{O}((T-T^{\prime})T^{\prime 2})$. Since $T^{\prime}$ increases gradually during the day and remains moderate in practical settings, the update remains computationally tractable. Moreover, conditional parameters are computed once per component and cached during sampling, and sparsity in mixture weights further reduces effective run time.
@@ IV-B3 Quantile-based
We use the same ensembles to form quantile forecasts $y_{t,n,i}^{T^{\prime}:}$ representing the $q_{i}$-th quantile of the ensemble set $\{x_{t,n,s}^{T^{\prime}:}\}_{s=1}^{S}$. Here, $\{q_{i}\}_{i=1}^{Q}\subset[0,1]^{Q}$ is the collection of $Q$ quantile levels with $q_{i}<q_{i+1},~\forall i$. After obtaining the quantile forecasts for each test day, the pinball loss function is used as an approximation of the well-known continuous ranked probability score (CRPS): [6]
@@ IV-C Baselines
The performance of the proposed intraday forecaster is heavily tied to the modelling accuracy of the day-ahead forecasting distribution. To isolate the effects of the updating mechanism on the day-ahead forecaster, we define a baseline model and a synthetic test set. Note that the purpose of this baseline is to isolate the effect of the Bayesian updating operator while keeping the day-ahead distribution fixed, rather than to compare performance with alternative retraining-based forecasting pipelines.
@@ IV-D Tuning the number of GMM components
IV-E Implementation Details
The CVAE encoder and decoder were implemented as fully connected feed-forward networks with three hidden layers of 1000 neurons each, using ReLU activations. Dropout with probability 0.2 and batch normalisation were applied after each hidden layer. The latent dimensions were set equal to the profile length $L=T$. The pattern dictionary size in the PDCC covariance model is chosen heuristically as $V=100$ for the Gipuzkoa SM dataset and $V=200$ for the Utrecht PV dataset. The models were trained using the Adam optimiser with default parameters in PyTorch 2.4.0 and a batch size of 1024. An early stopping mechanism and a learning rate scheduler were employed based on validation performance. Further implementation details can be found in [2].
@@ V Results and Discussion
Finally, we compare performance over real and synthetic test sets in Fig. 6 and Fig. 7. For the Gipuzkoa SM dataset, performance traces are largely consistent with each other, except for the NLL metric, which suggests that the real test set contains more typical, and therefore easier-to-predict, samples. For the Utrecht PV dataset, on the other hand, the model performs better on the synthetic test set, as expected. Despite this discrepancy, we observe that the update mechanism still provides a performance gain on the real test set in all metrics. The comparatively modest improvement observed for real PV data is likely attributable to the dominant influence of exogenous weather forecast uncertainty and rapid cloud dynamics, which limit the predictive value of early-day measurements for later hours. In addition, the day-ahead model already captures much of the deterministic diurnal solar generation pattern, leaving less room for improvement through conditioning. Consequently, intraday updates yield smaller but still consistent gains across metrics.
@@ VI Conclusion
The experimental results on both electricity consumption and PV generation datasets confirm that the updating mechanism consistently improves forecast accuracy across likelihood-, sample-, quantile-, and point-based metrics. The improvements are most noticeable for time steps with strong temporal correlation to the observed data, including correlations that extend over longer ranges. This highlights the method’s ability to effectively incorporate real-time information. Moreover, the use of pattern dictionary–based covariance structures enhances both the baseline modelling capability and the quality of intraday updates, indicating that intraday conditioning directly benefits from more accurate joint covariance estimation.